# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v2

@@ Abstract
The performance of a model trained with noisy labels is often improved by simply retraining the model with its own predicted hard labels (i.e., $1$/00 labels). Yet, a detailed theoretical characterization of this phenomenon is lacking. In this paper, we theoretically analyze retraining in a linearly separable setting with randomly corrupted labels given to us and prove that retraining can improve the population accuracy obtained by initially training with the given (noisy) labels. To the best of our knowledge, this is the first such theoretical result. Retraining finds application in improving training with local label differential privacy (DP) which involves training with noisy labels. We empirically show that retraining selectively on the samples for which the predicted label matches the given label significantly improves label DP training at no extra privacy cost; we call this consensus-based retraining. As an example, when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining.
@@ 1 Introduction
The motivation for consensus-based retraining is that matching the predicted and given labels can potentially yield a smaller but much more accurate subset compared to the entire set; such a filtering effect can further improve the model’s performance. As we show in Section 5 (see Tables 3 and 7), this intuition bears out in practice.
@@
Figure 1: Retraining Intuition. Samples to the right (respectively, left) of the separator (black vertical line in the middle) and colored blue (respectively, red) have actual label $+1$ (respectively, $-1$). For both classes, the incorrectly labeled samples are marked by crosses ($\times$), whereas the correctly labeled samples are marked by dots ($\circ$) of the appropriate color. The amount of label noise and the number of training samples are the same in 1(a) and 1(b). The top and bottom plots show the joint scatter plot of the training samples with the (noisy) labels given to us and the labels predicted by the model after training with the given labels, respectively. Notice that in 1(a), the model correctly predicts the labels of several samples that were given to it with the wrong label – especially, those that are far away from the separator. This is not quite the case in 1(b). This difference gets reflected in the performance on the test set after retraining. Specifically, in 1(a), retraining increases the test accuracy to $97.67\%$ from $89\%$. However, retraining yields no improvement in 1(b). So the success of retraining depends on the degree of separation between the classes.
@@
There are plenty of ideas revolving around training a model with its own predictions, the two most common ones being self-training (Scudder,, 1965; Yarowsky,, 1995; Lee et al.,, 2013) and self-distillation (Furlanello et al.,, 2018; Mobahi et al.,, 2020); we discuss these and important differences from retraining in Section 2. However, from a theoretical perspective, we are not aware of any work proving that retraining a model with its predicted hard labels can be beneficial in the presence of label noise in any setting. In Section 4, we derive the first theoretical result (to our knowledge) showing that full retraining with hard labels improves model accuracy.
@@
The primary reason for our interest in retraining is that it turned out to be a simple yet effective way to improve training with local [^1] In this paper, we focus only on local label DP. So throughout the paper, we will mostly omit the word “local” for conciseness. label differential privacy (DP) whose goal is to safeguard the privacy of labels in a supervised ML problem by injecting label noise (see Section 3 for a formal definition). Label DP is used in scenarios where only the labels are considered sensitive, e.g., advertising, recommendation systems, etc. (Ghazi et al.,, 2021). Importantly, retraining can be applied on top of any label DP training algorithm at no extra privacy cost. Our main algorithmic contribution is empirically demonstrating the efficacy of consensus-based retraining in improving label DP training (Section 5). Three things are worth clarifying here. First, as a meta-idea, retraining is not particularly new; however, its application – especially with consensus-based filtering – as a light-weight way to improve label DP training at no extra privacy cost is new to our knowledge. Second, we are not advocating consensus-based retraining as a SOTA general-purpose algorithm for learning with noisy labels. Third, we do not view full retraining to be an algorithmic contribution; we consider it for theoretical analysis and as a baseline for consensus-based retraining.
@@
  • In Section 4, we consider a linearly separable binary classification problem wherein the data (feature) dimension is $d$, and we are given randomly flipped labels with the label flipping probability being $p<\frac{1}{2}$ independently for each sample. Our main result is proving that full retraining with the predicted hard labels improves the population accuracy obtained by initially training with the given labels, provided that $p$ is close enough to $\frac{1}{2}$ and the dataset size $n$ satisfies $\frac{d\log n}{(1-2p)}\lesssim n\lesssim\frac{d^{2}}{(1-2p)^{2}}$ ($\lesssim$” means “bounded asymptotically”, ignoring constant factor multiples); see Remark 4.10 for details. In addition, our results show that retraining becomes more beneficial as the amount of label noise (i.e., $p$) increases or as the degree of separation between the classes increases. To the best of our knowledge, these are the first theoretical results quantifying the benefits of retraining with predicted hard labels in the presence of label noise. The analysis of retraining is particularly challenging because of the dependence of the predicted labels on the entire training set and the non-uniform/sample-dependent nature of label noise in the predicted labels; see the discussion after the statement of Theorem 4.8.
  • In Section 5, we show the promise of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) as a simple way to improve the performance of any label DP algorithm, at no extra privacy cost. As an example, when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining (see Table 2). The corresponding improvement for a small BERT model trained on AG News Subset (a news classification dataset) with $\epsilon=0.5$ label DP is $11.7\%$ (see Table 6).
@@ 2 Related Work
Self-Training (ST). Retraining is similar in spirit to ST (Scudder,, 1965; Yarowsky,, 1995; Lee et al.,, 2013; Sohn et al.,, 2020) which is the process of progressively training a model with its own predicted hard labels in the semi-supervised setting. Our focus in this work is on the fully supervised setting. This is different from ST (in the semi-supervised setting) which typically selects samples based on the model’s confidence and hence, we call our algorithmic idea of interest retraining to distinguish it from ST. In fact, we show that our consensus-based sample selection strategy leads to better performance than confidence-based sample selection in Appendix H. There is a vast body of work on ST and related ideas; see Amini et al., (2022) for a survey. On the theoretical side also, there are several papers showing and quantifying different kinds of benefits of ST and related ideas (Carmon et al.,, 2019; Raghunathan et al.,, 2020; Kumar et al.,, 2020; Chen et al.,, 2020; Oymak and Gulcu,, 2020; Wei et al.,, 2020; Zhang et al.,, 2022). But none of these works characterize the pros/cons of ST or any related algorithm in the presence of noisy labels. In contrast, we show that retraining can provably improve accuracy in the presence of label noise in Section 4. Empirically, ST-based ideas have been proposed to improve learning with noisy labels (Reed et al.,, 2014; Tanaka et al.,, 2018; Han et al.,, 2019; Nguyen et al.,, 2019; Li et al.,, 2020; Goel et al.,, 2022); but these works do not have rigorous theory. In the context of theory on label noise and model’s confidence, Zheng et al., (2020) show that if the model’s predicted score for the observed label is small, then the observed label is likely not equal to the true label. Note that the results of Zheng et al., (2020) pertain to the correctness of the observed labels, whereas our results pertain to the correctness of the predicted labels.
@@
Self-Distillation (SD). Retraining is also similar in principle to SD (Furlanello et al.,, 2018; Mobahi et al.,, 2020); the major difference is that soft labels (i.e., predicted raw probabilities) are used in SD, whereas we use hard labels in retraining. Specifically, in SD, a teacher model is first trained with provided hard labels and then its predicted soft labels are used to train a student model with the same architecture as the teacher. SD is usually employed with a temperature parameter (Hinton et al.,, 2015) to force the teacher and student models to be different; we do not have any such parameter in retraining as it uses hard labels. SD is known to ameliorate learning in the presence of noisy labels (Li et al.,, 2017) and this has been theoretically analyzed by Dong et al., (2019); Das and Sanghavi, (2023). Dong et al., (2019) propose their own SD algorithm that uses dynamically updated soft labels and provide some complicated conditions of when their algorithm can learn the correct labels in the presence of noisy labels. In contrast, we analyze retraining with fixed hard labels. Das and Sanghavi, (2023) analyze the standard SD algorithm in the presence of noisy labels with fixed soft labels but their analysis in the classification setting requires some strong assumptions such as access to the population, feature maps of all points in the same class having the same inner product, etc. We do not require such strong assumptions in this paper (in fact, we present sample complexity bounds). Moreover, Das and Sanghavi, (2023) have extra $\ell_{2}$-regularization in their objective function to force the teacher and student models to be different. We do not apply any extra regularization for retraining.
@@
Label Differential Privacy (DP). Label DP (described in detail in Section 3) is a relaxation of full-data DP wherein the privacy of only the labels (and not the features) is safeguarded (Chaudhuri and Hsu,, 2011; Beimel et al.,, 2013; Wang and Xu,, 2019; Ghazi et al.,, 2021; Malek Esmaeili et al.,, 2021; Ghazi et al.,, 2022; Badanidiyuru et al.,, 2023). In this work, we are not trying to propose a SOTA label DP algorithm (with an ingenious noise-injection scheme); instead, we advocate retraining as a simple post-processing step that can be applied on top of any label DP algorithm (regardless of the noise-injection scheme) to improve its performance, at no extra privacy cost. Similar to our goal, Tang et al., (2022) apply techniques from unsupervised and semi-supervised learning to improve label DP training. In particular, one of their steps involves keeping the given noisy label of a sample only if it matches a pseudo-label generated by unsupervised learning. This is similar in spirit to our consensus-based retraining scheme but a crucial difference is that we do not perform any unsupervised learning; we show that matching the given noisy label to the model’s own predicted label is itself pretty effective. Further, unlike our work, Tang et al., (2022) do not have any rigorous theory.
@@ 3 Preliminaries
Notation. For two functions $f(n)$ and $g(n)$, we write $f(n)\lesssim g(n)$ if there exists $n_{0}$ and a constant $C>0$ such that for all $n\geq n_{0}$, $f(n)\leq Cg(n)$. We also write $f(n)\ll g(n)$ to indicate that $f$ is asymptotically dominated by $g$, namely, for every fixed real number $C>0$, there exists $n(C)$ such that $g(n)\geq Cf(n)$ for all $n>n(C)$. For any positive integer $m\geq 1$, we denote the set $\{1,\ldots,m\}$ by $[m]$. Let $\bm{e}_{i}$ denote the $i^{\text{th}}$ canonical vector, namely, the vector of all zeros except a one in the $i^{\text{th}}$ coordinate. We denote the $\ell_{2}$ norm of a vector $\bm{v}$ by $\left\|\bm{v}\right\|_{\ell_{2}}$, and the operator norm of a matrix ${\bm{M}}$ by $\left\|{\bm{M}}\right\|$. The unit $d$-dimensional sphere (i.e., the set of $d$-dimensional vectors with unit norm) is denoted by $\mathbb{S}^{d-1}$. For a random variable $X$, its sub-gaussian norm, denoted by $\|X\|_{\psi_{2}}$, is defined as $\|X\|_{\psi_{2}}:=\sup_{q\geq 1}q^{-1/2}\big{(}\operatorname{\mathbb{E}}|X|^{q}\big{)}^{1/q}$. [^2] We refer to (Vershynin,, 2010, Lemma 5.5) for other equivalent definitions. In addition, for random vector ${\bm{X}}\in\mathbb{R}^{d}$, its sub-gaussian norm is defined as $\|{\bm{X}}\|_{\psi_{2}}=\sup_{{\bm{z}}\in\mathbb{S}^{d-1}}\|\langle{\bm{X}},{\bm{z}}\rangle\|_{\psi_{2}}$. We denote the CDF and complementary CDF (CCDF) of a standard normal variable (i.e., distributed as ${\mathcal{N}}(0,1)$) by $\Phi(.)$ and $\Phi^{\text{c}}(.)$, respectively.
Definition 3.1 (Label Differential Privacy (DP))
@@
Local label DP training involves injecting noise into the labels and then training with these noisy labels. The simplest way of injecting label noise to ensure label DP is randomized response (RR) introduced by Warner, (1965). Specifically, suppose we require $\epsilon$-labelDP for a problem with $C$ classes, then the distribution of the output $\widehat{y}$ of RR when the true label is $y$ is as follows:
$\mathbb{P}\big{(}\widehat{y}=z\big{)}=\begin{cases}\frac{e^{\epsilon}}{e^{\epsilon}+C-1}\quad\text{ for }z=y,\\ \frac{1}{e^{\epsilon}+C-1}\quad\text{ otherwise}.\end{cases}$
Our label noise model in Section 4 (eq. 4.2) is actually RR for 2 classes. Based on vanilla RR, more sophisticated ways to inject label noise for better performance under label DP have been proposed (Ghazi et al.,, 2021; Malek Esmaeili et al.,, 2021). Our empirical results in Section 5 are with RR and the method of Ghazi et al., (2021).
@@ 4 Full Retraining in the Presence of Label Noise: Theoretical Analysis
Problem Setting. We consider binary classification under a discriminative mixtures of Gaussian data model with a positive margin. We will first describe the classical Gaussian mixture model and then propose a new model which is endowed with a positive margin.
In the classical Gaussian mixture model, each data point belongs to one of two classes {±$\{\pm 1\}$ with corresponding probabilities $\pi_{+}$, $\pi_{-}$, such that $\pi_{+}+\pi_{-}=1$. Denoting by $y_{i}\in\{-1,+1\}$ the label for data point $i$, the feature vectors $\bm{x}_{i}\in\mathbb{R}^{d}$, for $i\in[n]$, are generated independently as $\bm{x}_{i}\sim{\mathcal{N}}(y_{i}{\bm{{\mu}}},\bm{\Sigma})$, where ${\bm{{\mu}}}\in\mathbb{R}^{d}$ and $\bm{\Sigma}\in\mathbb{R}^{d\times d}$. In other words the mean of feature vectors are ±$\pm{\bm{{\mu}}}$ depending on its class, and $\bm{\Sigma}$ is the covariance of the features. We let $\gamma:=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}$.
Gaussian mixture model with positive margin: In this model, each data point $(\bm{x},y)$ is generated independently by first sampling the label $y\in\{-1,+1\}$ with the corresponding probabilities $\pi_{-},\pi_{+}$, and then generating the feature vector $\bm{x}\in\mathbb{R}^{d}$ as:
@@
  • $u$ is drawn independently for each sample from a common sub-gaussian distribution with positive support ($u>0$).
@@
  • $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ is the covariance matrix in the space orthogonal to ${\bm{{\mu}}}$. We have $\bm{\Sigma}^{1/2}{\bm{{\mu}}}=\bm{0}$ and $\bm{\Sigma}$ is of rank $d-1$.
We let $\bm{\Sigma}$ be an arbitrary positive semi-definite matrix to handle general covariances. Let $\lambda_{\min}>0$ and $\lambda_{\max}$ denote the minimum non-zero eigenvalue and the maximum eigenvalue of $\bm{\Sigma}$, respectively. Also without loss of generality, we assume that $u$ has unit sub-gaussian norm ($\|u\|_{\psi_{2}}=1$).
Note that under this model, the projection of datapoints on the orthogonal space of ${\bm{{\mu}}}$ is distributed as normal vectors. On the direction ${\bm{{\mu}}}$, we have $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}$, and so the randomness of data along this direction is modeled in $u$. Since $u$ has a positive support, $\text{sign}(\langle\bm{x},{\bm{{\mu}}}\rangle)=y$ and therefore ${\bm{{\mu}}}$ is a separator of the data (based on its labels). In addition, we have a margin of at least $\gamma=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}$; this is because $\frac{y\langle\bm{x},{\bm{{\mu}}}\rangle}{\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}}=(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}\geq\gamma$. We will mostly refer to the margin $\gamma$ as the “degree of separation”.
@@
4.1 Vanilla Training
Given the training set $\mathcal{T}=\{(\bm{x}_{i},\widehat{y}_{i})\}_{i\in[n]}$, we consider the following linear classifier (Carmon et al.,, 2019):

### v3

@@ Abstract
The performance of a model trained with noisy labels is often improved by simply retraining the model with its own predicted hard labels (i.e., $1$/00 labels). Yet, a detailed theoretical characterization of this phenomenon is lacking. In this paper, we theoretically analyze retraining in a linearly separable binary classification setting with randomly corrupted labels given to us and prove that retraining can improve the population accuracy obtained by initially training with the given (noisy) labels. To the best of our knowledge, this is the first such theoretical result. Retraining finds application in improving training with local label differential privacy (DP) which involves training with noisy labels. We empirically show that retraining selectively on the samples for which the predicted label matches the given label significantly improves label DP training at no extra privacy cost; we call this consensus-based retraining. As an example, when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain more than $6\%$ improvement in accuracy with consensus-based retraining.
Machine Learning, ICML
@@ 1 Introduction
The motivation for consensus-based retraining is that matching the predicted and given labels can potentially yield a smaller but much more accurate subset compared to the entire set; such a filtering effect can further improve the model’s performance. As we show in Section 5 (see Tables 3 and 11), this intuition bears out in practice.
@@
Figure 1: Retraining Intuition. Samples to the right (respectively, left) of the separator (black vertical line in the middle) and colored blue (respectively, red) have actual label $+1$ (respectively, $-1$). For both classes, the incorrectly labeled samples are marked by crosses ($\times$), whereas the correctly labeled samples are marked by dots ($\circ$) of the appropriate color. The amount of label noise and the number of training samples are the same in 1(a) and 1(b). The top and bottom plots show the joint scatter plot of the training samples with the (noisy) labels given to us and the labels predicted by the model after training with the given labels, respectively. Notice that in 1(a), the model correctly predicts the labels of several samples that were given to it with the wrong label – especially, those that are far away from the separator. This is not quite the case in 1(b). This difference gets reflected in the performance on the test set after retraining. Specifically, in 1(a), retraining increases the test accuracy to $97.67\%$ from $89\%$. However, retraining yields no improvement in 1(b). So the success of retraining depends on the inter-class separation; in particular, retraining is beneficial when the classes are well-separated.
@@
There are plenty of ideas revolving around training a model with its own predictions, the two most common ones being self-training (Scudder, 1965; Yarowsky, 1995; Lee et al., 2013) and self-distillation (Furlanello et al., 2018; Mobahi et al., 2020); we discuss these and important differences from retraining in Section 2. However, from a theoretical perspective, we are not aware of any work proving that retraining a model with its predicted hard labels can be beneficial in the presence of label noise in any setting. In Section 4, we derive the first theoretical result (to our knowledge) showing that full retraining with hard labels improves model accuracy.
@@
The primary reason for our interest in retraining is that it turned out to be a simple yet effective way to improve training with local [^1] In this paper, we focus only on local label DP. So throughout the paper, we will mostly omit the word “local” for conciseness. label differential privacy (DP) whose goal is to safeguard the privacy of labels in a supervised ML problem by injecting label noise (see Section 3 for a formal definition). Label DP is used in scenarios where only the labels are considered sensitive, e.g., advertising, recommendation systems, etc. (Ghazi et al., 2021). Importantly, retraining can be applied on top of any label DP training algorithm at no extra privacy cost. Our main algorithmic contribution is empirically demonstrating the efficacy of consensus-based retraining in improving label DP training (Section 5). Three things are worth clarifying here. First, as a meta-idea, retraining is not particularly new; however, its application – especially with consensus-based filtering – as a light-weight way to improve label DP training at no extra privacy cost is new to our knowledge. Second, we are not advocating consensus-based retraining as a SOTA general-purpose algorithm for learning with noisy labels. Third, we do not view full retraining to be an algorithmic contribution; we consider it for theoretical analysis and as a baseline for consensus-based retraining.
@@
(a) In Section 4, we consider a linearly separable binary classification problem wherein the data (feature) dimension is $d$, and we are given randomly flipped labels with the label flipping probability being $p<\frac{1}{2}$ independently for each sample. Our main result is proving that full retraining with the predicted hard labels improves the population accuracy obtained by initially training with the given labels, provided that $p$ is close enough to $\frac{1}{2}$ and the dataset size $n$ satisfies $\frac{d}{(1-2p)^{2}}\log\frac{d}{(1-2p)^{2}}\lesssim n\lesssim\frac{d^{2}}{(1-2p)^{2}}$ ($\lesssim$” means bounded asymptotically, ignoring constant factor multiples); see Remark 4.10 for details. In addition, our results show that retraining becomes more beneficial as the amount of label noise (i.e., $p$) increases or as the degree of separation between the classes increases. To the best of our knowledge, these are the first theoretical results quantifying the benefits of retraining with predicted hard labels in the presence of label noise. The analysis of retraining is particularly challenging because of the dependence of the predicted labels on the entire training set and the non-uniform/sample-dependent nature of label noise in the predicted labels; see the discussion after the statement of Theorem 4.8.
(b) In Section 5, we show the promise of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) as a simple way to improve the performance of any label DP algorithm, at no extra privacy cost. For e.g., when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining (see Table 2). The corresponding improvement for a small BERT model trained on AG News Subset (a news classification dataset) with $\epsilon=0.5$ label DP is $11.7\%$ (see Table 10).
@@ 2 Related Work
Self-training (ST). Retraining is similar in spirit to ST (Scudder, 1965; Yarowsky, 1995; Lee et al., 2013; Sohn et al., 2020) which is the process of progressively training a model with its own predicted hard labels in the semi-supervised setting. Our focus in this work is on the fully supervised setting. This is different from ST (in the semi-supervised setting) which typically selects samples based on the model’s confidence and hence, we call our algorithmic idea of interest retraining to distinguish it from ST. In fact, we show that our consensus-based sample selection strategy leads to better performance than confidence-based sample selection in Appendix J. There is a vast body of work on ST and related ideas; see Amini et al. (2022) for a survey. On the theoretical side also, there are several papers showing and quantifying different kinds of benefits of ST and related ideas (Carmon et al., 2019; Raghunathan et al., 2020; Kumar et al., 2020; Chen et al., 2020; Oymak & Gulcu, 2020; Wei et al., 2020; Zhang et al., 2022). But none of these works characterize the pros/cons of ST or any related algorithm in the presence of noisy labels. In contrast, we show that retraining can provably improve accuracy in the presence of label noise in Section 4. Empirically, ST-based ideas have been proposed to improve learning with noisy labels (Reed et al., 2014; Tanaka et al., 2018; Han et al., 2019; Nguyen et al., 2019; Li et al., 2020; Goel et al., 2022); but these works do not have rigorous theory. In the context of theory on label noise and model’s confidence, Zheng et al. (2020) show that if the model’s predicted score for the observed label is small, then the observed label is likely not equal to the true label. Note that the results of Zheng et al. (2020) pertain to the correctness of the observed labels, whereas our results pertain to the correctness of the predicted labels.
@@
Self-distillation (SD). Retraining is also similar in principle to SD (Furlanello et al., 2018; Mobahi et al., 2020), where a teacher model is first trained with provided hard labels and then its predicted soft labels are used to train a student model with the same architecture as the teacher. In contrast, we use predicted hard labels in retraining. Also, SD is usually employed with a temperature parameter (Hinton et al., 2015) to force the teacher and student models to be different; we do not have any such parameter in retraining. SD is known to ameliorate learning in the presence of noisy labels (Li et al., 2017) and this has been theoretically analyzed by Dong et al. (2019); Das & Sanghavi (2023). Dong et al. (2019) propose their own SD algorithm that uses dynamically updated soft labels and provide some complicated conditions of when their algorithm can learn the correct labels in the presence of noisy labels. In contrast, we analyze retraining with fixed hard labels. Das & Sanghavi (2023) analyze the standard SD algorithm in the presence of noisy labels with fixed soft labels but their analysis in the classification setting requires some strong assumptions such as access to the population, feature maps of all points in the same class having the same inner product, etc. We do not require such strong assumptions in this paper (in fact, we present sample complexity bounds).
@@
Label differential privacy (DP). Label DP (described in detail in Section 3) is a relaxation of full-data DP wherein the privacy of only the labels (and not the features) is safeguarded (Chaudhuri & Hsu, 2011; Beimel et al., 2013; Wang & Xu, 2019; Ghazi et al., 2021; Malek Esmaeili et al., 2021; Ghazi et al., 2022; Badanidiyuru et al., 2023). In this work, we are not trying to propose a SOTA label DP algorithm (with an ingenious noise-injection scheme); instead, we advocate retraining as a simple post-processing step that can be applied on top of any label DP algorithm (regardless of the noise-injection scheme) to improve its performance, at no extra privacy cost. Similar to our goal, Tang et al. (2022) apply techniques from unsupervised and semi-supervised learning to improve label DP training. In particular, one of their steps involves keeping the given noisy label of a sample only if it matches a pseudo-label generated by unsupervised learning. This is similar in spirit to our consensus-based retraining scheme but a crucial difference is that we do not perform any unsupervised learning; we show that matching the given noisy label to the model’s own predicted label is itself pretty effective. Further, unlike our work, Tang et al. (2022) do not have any rigorous theory.
@@ 3 Preliminaries
Notation. For two functions $f(n)$ and $g(n)$, we write $f(n)\lesssim g(n)$ if there exists $n_{0}$ and a constant $C>0$ such that for all $n\geq n_{0}$, $f(n)\leq Cg(n)$. For any positive integer $m\geq 1$, we denote the set $\{1,\ldots,m\}$ by $[m]$. Let $\bm{e}_{i}$ denote the $i^{\text{th}}$ canonical vector, namely, the vector of all zeros except a one in the $i^{\text{th}}$ coordinate. We denote the $\ell_{2}$ norm of a vector $\bm{v}$ by $\left\|\bm{v}\right\|_{\ell_{2}}$, and the operator norm of a matrix ${\bm{M}}$ by $\left\|{\bm{M}}\right\|$. The unit $d$-dimensional sphere (i.e., the set of $d$-dimensional vectors with unit norm) is denoted by $\mathbb{S}^{d-1}$. For a random variable $X$, its sub-gaussian norm, denoted by $\|X\|_{\psi_{2}}$, is defined as $\|X\|_{\psi_{2}}:=\sup_{q\geq 1}q^{-1/2}\big{(}\operatorname{\mathbb{E}}|X|^{q}\big{)}^{1/q}$. [^2] Refer to Vershynin (2010) for other equivalent definitions. In addition, for random vector ${\bm{X}}\in\mathbb{R}^{d}$, its sub-gaussian norm is defined as $\|{\bm{X}}\|_{\psi_{2}}=\sup_{{\bm{z}}\in\mathbb{S}^{d-1}}\|\langle{\bm{X}},{\bm{z}}\rangle\|_{\psi_{2}}$. We denote the CDF and complementary CDF (CCDF) of a standard normal variable (i.e., distributed as ${\mathcal{N}}(0,1)$) by $\Phi(.)$ and $\Phi^{\text{c}}(.)$, respectively.
Definition 3.1 (Label differential privacy (DP)).
@@
Local label DP training involves injecting noise into the labels and then training with these noisy labels. The simplest way of injecting label noise to ensure label DP is randomized response (RR) introduced by Warner (1965). Specifically, suppose we require $\epsilon$-labelDP for a problem with $C$ classes, then the distribution of the output $\widehat{y}$ of RR when $y$ is the true label is given by: $\mathbb{P}\left(\widehat{y}=y\right)=\frac{e^{\epsilon}}{e^{\epsilon}+C-1}$, and $\mathbb{P}\left(\widehat{y}=z\right)=\frac{1}{e^{\epsilon}+C-1}$ for any $z\neq y$. Our label noise model in Section 4 (Equation 4.2) is actually RR for 2 classes. Based on vanilla RR, more sophisticated ways to inject label noise for better performance under label DP have been proposed (Ghazi et al., 2021; Malek Esmaeili et al., 2021). Our empirical results in Section 5 are with RR and the method of Ghazi et al. (2021).
@@ 4 Full Retraining in the Presence of Label Noise: Theoretical Analysis
Problem setting: We consider binary classification in a linearly separable [^3] More specifically, there exists a linear separator that perfectly classifies the data points according to their labels. setting similar to the classical Gaussian mixture model. We will first describe the classical Gaussian mixture model and then describe our setting of interest, which is endowed with a positive margin.
In the classical Gaussian mixture model, each data point belongs to one of two classes $\{+1,-1\}$ with corresponding probabilities $\pi_{+}$, $\pi_{-}$, such that $\pi_{+}+\pi_{-}=1$. The feature vector $\bm{x}\in\mathbb{R}^{d}$ for a data point with label $y\in\{+1,-1\}$ is drawn independently from ${\mathcal{N}}(y{\bm{{\mu}}},\bm{\Sigma})$, where ${\bm{{\mu}}}\in\mathbb{R}^{d}$ and $\bm{\Sigma}\in\mathbb{R}^{d\times d}$. In other words the mean of the feature vector is ±$\pm{\bm{{\mu}}}$ depending on the data point’s label, and $\bm{\Sigma}$ is the covariance matrix of the feature vector. We let $\gamma:=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}$.
Gaussian mixture model with positive margin: Here each data point $(\bm{x},y)$ is generated independently by first sampling the label $y\in\{+1,-1\}$ with probabilities $\pi_{+},\pi_{-}$, and then generating the feature vector $\bm{x}\in\mathbb{R}^{d}$ as:
@@
  • $u>0$ is drawn from a sub-gaussian distribution with unit sub-gaussian norm (i.e., $\|u\|_{\psi_{2}}=1$).
@@
  • $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ is the covariance matrix in the space orthogonal to ${\bm{{\mu}}}$. We have $\bm{\Sigma}^{1/2}{\bm{{\mu}}}=\bm{0}$ and $\bm{\Sigma}$ is of rank $d-1$. Also, $\lambda_{\min}>0$ and $\lambda_{\max}$ are the minimum non-zero eigenvalue and the maximum eigenvalue of $\bm{\Sigma}$.
Note that under this data model, the projection of datapoints on the space orthogonal to ${\bm{{\mu}}}$ is Gaussian. Along ${\bm{{\mu}}}$, we have $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}$, and so the randomness of data along ${\bm{{\mu}}}$ is due to $u$. Further, since $u>0$, $\text{sign}(\langle\bm{x},{\bm{{\mu}}}\rangle)=y$. Thus, ${\bm{{\mu}}}$ is a separator for the data w.r.t. the labels. In addition, we have a margin of at least $\gamma=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}$; this is because $\frac{y\langle\bm{x},{\bm{{\mu}}}\rangle}{\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}}=(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}\geq\gamma$. We will mostly refer to the margin $\gamma$ as the “degree of separation”.
@@
4.1 Initial Training
Given the training set $\mathcal{T}=\{(\bm{x}_{i},\widehat{y}_{i})\}_{i\in[n]}$, we consider the following linear classifier model [^5] This is a simplification to the least squares’ solution (LSS) obtained by excluding the empirical covariance matrix’s inverse. The LSS can be analyzed by bounding the deviation of the empirical covariance matrix from the population covariance matrix (which shrinks as $n\to\infty$), then adapting our current analysis to features pre-multiplied by the population covariance matrix’s inverse, and accounting for this deviation. This would only make the math more tedious without adding any meaningful insights. (also used in Carmon et al. (2019)):